OpenStudy (jessicaloraine):

I would like to know how to rewrite this expression with common logarithms with log base 2 to the 10

OpenStudy (unklerhaukus):

use the change of base formula \[\log_ba=\frac{\log_ca}{\log_cb}\]

OpenStudy (unklerhaukus):

i.e. \[\log x = \log_{10}x=\frac{\log_2x}{\log_210}=\frac{\operatorname{lb}x}{\operatorname{lb}10}\]or \[\ln x = \log_ex=\frac{\log_2x}{\log_2e}=\frac{\operatorname{lb}x}{\operatorname{lb}e}\]

OpenStudy (jessicaloraine):

thank you very much

OpenStudy (unklerhaukus):

\[\ddot\smile \]